Bill makes car payments of $273/month for the next 3 years.


His car loan has an interest rate of 2.6%, discounted monthly.


What was the initial price of the car?



1
Expert's answer
2021-05-02T08:12:48-0400

calculating the initial price of the car

"Pv=PMT*\frac{1-(1+r)^{-n}}{r}"

where Pv is the present(initial value)

PMT is the amount of monthly payment

r is the rate of discount

n is the number of repayment periods.

we need to find the number of periods in months; 3*12=36 months

"Pv=273*\frac{1-(1+0.026)^{-36}}{0.026}"

=6332.41